# On the modified $J$-equation

### v3

@@ 4.1.2. Gluing argument
The above discussion prepares the ground for constructing a subsolution on $X$. We are now ready to begin the proof.
@@
where we used Proposition 2.2 and the fact that $\mathrm{Re}(v)$ is tangent to $X$. Hence, by Proposition 2.15, if we set
@@ 5.3. Interior $C^{2}$-estimates
for some constant $\delta>0$. Adding an affine linear function to $g$ amounts to a translation of the moment polytope $P$. Since the equation (5.7) and the condition (5.8) remain unchanged under this operation, we may assume that $g(0)=0$ and $\nabla g(0)=0$. In particular, since $0\in P$ we obtain:
@@
with constant coefficients $\widetilde{a}_{\infty,ij}$, $\widetilde{b}_{\infty}$ due to (5.15). Since $L_{\infty}(y_{0,\infty})\geqslant\delta>0$, after a linear change of coordinates and scaling of $\widetilde{h}_{\infty}$ we can apply the Liouville rigidity result [CS17, Lemma 31] to conclude that $\widetilde{h}_{\infty}$ is a quadratic polynomial. This contradicts the fact that $|D^{3}\widetilde{h}_{\infty}(0)|=1$, completing the proof. ∎
@@ 6. A numerical criterion for extremal Kähler metrics
In this section, we give a proof of Theorem 1.7 by adapting Theorem 1.6 to the arguments in [DJ25] [^3] More precisely, Delcroix–Jubert [DJ25] obtain an existence result for $(\mathrm{v},\mathrm{w})$-weighted cscK metrics in their context, and extremal Kähler metrics arises as the special case with $\mathrm{v}=1$ and $\mathrm{w}$ being the affine linear function correspoinding to the extremal vector field $v_{\rm ext}$.. First, let us consider a general compact complex manifold $X$ and a maximal compact torus $T\subset\operatorname{Aut}_{\rm red}(X)$. Let $\widehat{\chi}$ be a $T$-invariant Kähler form, and let $\beta$ denote its Kähler class. We choose the extremal vector field $v_{\rm ext}$ so that $\mathrm{Im}(v_{\rm ext})\in\mathfrak{t}$. The following discussion is entirely based on [DJ25]. However, since there are differences in conventions such as the definitions of functionals, we briefly summarize and unify them in our setting for the reader’s convenience. We define the modified $K$-energy
@@
This is well-known as the Chen–Tian formula (cf. [Che00, Tia97a, Tia97b]). Define the $T^{\mathbb{C}}$-reduced delta invariant of the Kähler class $\beta$ by
@@
Note that the value $\delta_{\beta}^{T^{\mathbb{C}}}$ does not depend on the choice of reference form $\widehat{\chi}\in\beta$, as can be verified by an argument similar to the proof of Theorem 1.3. By using Tian’s alpha invariant, one can also see that $\delta_{\beta}^{T^{\mathbb{C}}}>0$ [DJ25, Section 3.2]. In view of (6.1), the invariant $\delta_{\beta}^{T^{\mathbb{C}}}$ quantitatively measures the coercivity of the modified $K$-energy $K_{v_{\rm ext}}$. Let $\varepsilon>0$ be a constant. For each $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, let us take $\tau\in T^{\mathbb{C}}$ that minimizes $J^{\widehat{\chi}}(\phi_{\tau})$ so that $J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi_{\tau})=J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)=J^{\widehat{\chi}}(\phi_{\tau})$. As in [DJ25, Lemma 5.4], using the $T^{\mathbb{C}}$-invariance of $K_{v_{\rm ext}}$, Chen–Tian formula (6.1), and definition of $\delta_{\beta}^{T^{\mathbb{C}}}$, we obtain
@@
Thus, showing that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive reduces to the problem of establishing a lower bound for the functional $J_{-v_{\rm ext}}^{(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\widehat{\chi}-\operatorname{Ric}(\widehat{\chi})}$. Moreover, if the class $\alpha:=(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler, then we can choose a Kähler form $\omega\in\alpha$, and define a modified $J$-functional
@@
Therefore, by applying Theorem 1.6 to the equation (6.4) in the toric case, we know that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive under the assumptions in Theorem 1.7. Then it follows from [NJL25, Theorem 1.2] that the Kähler class $\beta$ admits a $T$-invariant extremal Kähler metric. This completes the proof of Theorem 1.7.

### v4

@@ 4.1.2. Gluing argument
The above discussion prepares the ground for constructing a subsolution on $X$. We now proceed to the proof.
@@
where we used Proposition 2.2 (2) and the fact that $\mathrm{Re}(v)$ is tangent to $X$. Hence, by Proposition 2.15, if we set
@@ 5.3. Interior $C^{2}$-estimates
for some constant $\delta>0$. Adding an affine linear function to $g$ amounts to a parallel translation of the moment polytope $P$. Since the equation (5.7) and the condition (5.8) remain unchanged under this operation, we may assume that $g(0)=0$ and $\nabla g(0)=0$. In particular, since $0\in P$ we obtain:
@@
with constant coefficients $\widetilde{a}_{\infty,ij}$, $\widetilde{b}_{\infty}$ due to (5.15). Since $L_{\infty}(y_{0,\infty})\geqslant\delta>0$, after a linear change of coordinates, we can apply the Liouville rigidity result [CS17, Lemma 31] to conclude that $\widetilde{h}_{\infty}$ is a quadratic polynomial. This contradicts the fact that $|D^{3}\widetilde{h}_{\infty}(0)|=1$, completing the proof. ∎
@@ 6. A numerical criterion for extremal Kähler metrics
Let $X$ be an $n$-dimensional compact Kähler manifold and $T\subset\operatorname{Aut}_{\rm red}(X)$ a maximal compact torus. Let $\widehat{\chi}$ be a $T$-invariant Kähler form, and let $\beta$ denote its Kähler class. We choose the extremal vector field $v_{\rm ext}$ so that $\mathrm{Im}(v_{\rm ext})\in\mathfrak{t}$. The following discussion is entirely based on [DJ25]. However, since there are differences in conventions such as the definitions of functionals, we briefly summarize and unify them in our setting for the reader’s convenience. We define the modified $K$-energy
@@
This is well-known as the Chen–Tian formula [Che00, Tia97a, Tia97b]. Define the $T^{\mathbb{C}}$-reduced delta invariant of the Kähler class $\beta$ by
@@
Note that the value $\delta_{\beta}^{T^{\mathbb{C}}}$ does not depend on the choice of reference form $\widehat{\chi}\in\beta$, as can be verified by an argument similar to the proof of Theorem 1.3. By using Tian’s alpha invariant, one can also see that $\delta_{\beta}^{T^{\mathbb{C}}}>0$ [DJ25, Section 3.2]. In view of (6.1), the invariant $\delta_{\beta}^{T^{\mathbb{C}}}$ quantitatively measures the coercivity of the modified $K$-energy $K_{v_{\rm ext}}$.
As an application to the existence problem of extremal Kähler metrics, we now prove the following theorem:
Theorem 6.1.
Let $X$ be an $n$-dimensional compact Kähler manifold, and $T$, $v_{\rm ext}$, $\widehat{\chi}$, $\beta$ as above. Assume that Conjecture 1.4 holds, and that there exists a constant $\varepsilon>0$ satisfying the following properties:
  1. (1)
  2. (2)
    For any $T$-invariant Kähler forms $\chi\in\beta$,
  3. (3)
    For any $T$-invariant $p$-dimensional irreducible subvarieties $Y\subset X$ with $p=1,\ldots,n-1$ and any $T$-invariant Kähler forms $\chi\in\beta$,
Proof.
For each $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, let us take $\tau\in T^{\mathbb{C}}$ that minimizes $J^{\widehat{\chi}}(\phi_{\tau})$ so that $J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi_{\tau})=J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)=J^{\widehat{\chi}}(\phi_{\tau})$. As in [DJ25, Lemma 5.4], using the $T^{\mathbb{C}}$-invariance of $K_{v_{\rm ext}}$, Chen–Tian formula (6.1), and definition of $\delta_{\beta}^{T^{\mathbb{C}}}$, we obtain
@@
Thus, showing that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive reduces to the problem of establishing a lower bound for the functional $J_{-v_{\rm ext}}^{(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\widehat{\chi}-\operatorname{Ric}(\widehat{\chi})}$. Since the class $\alpha:=(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler, we can choose a Kähler form $\omega\in\alpha$, and define the modified $J$-functional
@@
Therefore, assuming that Conjecture 1.4 is true, we can apply it to the equation (6.4), and deduce that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive under the assumptions in Theorem 6.1. Then it follows from [NJL25, Theorem 1.2] that the Kähler class $\beta$ admits a $T$-invariant extremal Kähler metric. This completes the proof of Theorem 6.1. ∎
We conclude this section with a few additional remarks. First, a variant of Theorem 6.1 was already obtained in [Che21, Corollary 1.5]. The formulation there is essentially the same as ours, except that we work in the $T^{\mathbb{C}}$-reduced setting.
We also emphasize that the above theorem is essentially established in [DJ25, Theorem 1.4] [^3] More precisely, Delcroix–Jubert [DJ25] obtain an existence result for $(\mathrm{v},\mathrm{w})$-weighted cscK metrics in their context, and extremal Kähler metrics arises as the special case with $\mathrm{v}=1$ and $\mathrm{w}$ being the affine linear function correspoinding to the extremal vector field $v_{\rm ext}$.. Our contribution lies in identifying a numerical characterization of the solvability of the modified $J$-equation (6.4), under the assumption that Conjecture 1.4 holds. In particular, the conjecture allows one to replace the analytic solvability condition by explicit intersection inequalities involving $T$-invariant subvarieties.
Moreover, in the toric case, Theorem 1.7 yields the toric analogue of Conjecture 1.4. In particular, a stronger statement holds: the conclusions remain valid even if $\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon$ in assumptions (1)–(3) of Theorem 6.1 is replaced by $\delta_{\beta}^{T^{\mathbb{C}}}$. This is due to the fact that, in (3), the objects under consideration are restricted to toric subvarieties, which form a finite set.